OpenStudy (anonymous):

What are the solutions of the equation? Graph and check the solutions. | x | + 10 = 1 A. x = 9 B. x = 9 or x = –9 C. x = 0 D. no solution

OpenStudy (anonymous):

You made a typo above ^. D is the correct answer.

OpenStudy (anonymous):

x+10=1 x=-9 or -x+10=1 -x=-9 x=9 However, when you plug in x as -9 or 9, the equation is false. Therefore, there is no solution. If you think about it, there is no way you can add 10 to a positive number and have the result be smaller than 10.